# The strong convergence phenomenon

### v1

@@ 4.3. Discussion: on the role of interpolation
In the special case that $X$ is a GUE matrix, we obtained a much better result than Lemma 4.7: the crossing identity (4.1) captures the effect of a crossing exactly. This identity can be iterated in the presence of multiple crossings, which results in the genus expansion for GUE matrices (see, e.g., [99, §1.7]). This is a rather special feature of classical random matrix models, however, and we do not know of any method that can meaningfully capture the effect of multiple crossings in the setting of arbitrarily structured random matrices.
@@ 5.1. Random lifts of graphs
We begin by recalling some basic notions that can be found, for example, in [71, §6].
@@ 5.1.1. Random lifts
From the perspective of the lower bound, there is nothing particularly special about $d$-regular graphs beside that they all have the same universal cover. Indeed, for any sequence of graphs with diverging number of vertices that have the same universal cover, the maximum nontrivial eigenvalue is asymptotically lower bounded by the spectral radius of the universal cover. This follows by a straightforward adaptation of Lemma 1.2, cf. [71, Theorem 6.6].
@@ 5.1.2. Old and new eigenvalues
Even though Corollary 5.2 is only concerned with the new eigenvalues of $G^{N}$, it implies the classical spectral gap property $\|A^{N}|_{1^{\perp}}\|\to\varrho$ whenever the base graph satisfies $\|A|_{1^{\perp}}\|\leq\varrho$. Another simple consequence is that whenever the base graph satisfies $\|A\|>\varrho$, the random lift $G^{N}$ is connected with probability $1-o(1)$; this holds if and only if $G$ has at least two cycles [73, Theorem 2].
@@ 5.2. Buser’s conjecture
Let $X$ be a hyperbolic surface, that is, a connected Riemannian surface of constant curvature $-1$. Then $X$ has the hyperbolic plane $\mathbb{H}$ as its universal cover, and we can in fact obtain $\𝑋\ΓX=\Gamma\backslash\mathbb{H}$ as a quotient of the hyperbolic plane by a Fuchsian group $\Gamma$ (i.e., a discrete subgroup of $\mathrm{PSL}_{2}(\mathbb{R})$) which is isomorphic to the fundamental group $\Gamma\simeq\pi_{1}(X)$.
@@ 5.2.1. Random covers
Let us begin with an informal discussion. The action of a Fuchsian group $\Gamma$ on $\mathbb{H}$ defines a Dirichlet fundamental domain $F$ whose translates $\{\gamma F:\gamma\in\Gamma\}$ tile $\mathbb{H}$; $F$ is a polygon whose sides are given by $F\cap\gamma F$ and $F\cap\gamma^{-1}F$ for some generating set $\gamma\in\{\gamma_{1},\ldots,\gamma_{s}\}$ of $\Gamma$. Then $\𝑋\ΓX=\Gamma\backslash\mathbb{H}$ is obtained from $F$ by gluing each pair of sides $F\cap\gamma_{i}F$ and $F\cap\gamma_{i}^{-1}F$. See Figure 5.2 and [12, Chapter 9].
$F$$\scriptstyle\gamma_{1}F$$\scriptstyle\gamma_{2}^{\text{-}1}F$$\scriptstyle\gamma_{1}^{\text{-}1}F$$\scriptstyle\gamma_{2}F$$\scriptstyle\gamma_{3}F$$\scriptstyle\gamma_{4}^{\text{-}1}F$$\scriptstyle\gamma_{3}^{\text{-}1}F$$\scriptstyle\gamma_{4}F$$\Γ\Gamma\backslash\mathbb{H}$
Figure 5.2. Illustration of a tiling of $\mathbb{H}$ (in the Poincaré disc model) by hyperbolic octagons; gluing the sides of the fundamental domain $F$ yields a genus $2$ surface.
@@
More formally, this construction can be implemented as follows. Fix a base surface $\𝑋\ΓX=\Gamma\backslash\mathbb{H}$ and a homomorphism $\pi_{N}\in\mathrm{Hom}(\Gamma,\mathbf{S}_{N})$. Define
@@
where we let $\gamma\in\Gamma$ act on ×$\mathbb{H}\times[N]$ as $\gamma(z,i)=(\gamma z,\pi_{N}(\gamma)i)$. Then $X^{N}$ is an $N$-fold cover of $X$, and every $N$-fold cover of $X$ arises in this manner for some choice of $\pi_{N}$; cf. [64, pp. 68–70] or [53, §14a and §16d].
@@
where $[g,h]=ghg^{-1}h^{-1}$. In this case, the random permutations $\sigma_{i}$ must be chosen to satisfy $[\sigma_{1},\sigma_{2}]\cdots[\sigma_{2g-1},\sigma_{2g}]=1$, which precludes them from being independent. The reason this issue does not arise for graphs is that the fundamental group of every graph is free, and thus there are no relations to be satisfied.
@@
    Finally, the most natural model of random covers of closed surfaces is to choose $\pi_{N}\in\mathrm{Hom}(\Gamma,\mathbf{S}_{N})$ uniformly at random, that is, choose $\sigma_{i}=\pi_{N}(\gamma_{i})$ uniformly at random among the set of tuples $\sigma_{1},\ldots,\sigma_{s}\in\mathbf{S}_{N}$ that satisfy the relation of $\Gamma$. This corresponds to choosing an $N$-fold cover of $X$ uniformly at random [91]. The challenge in analyzing this model is that $\sigma_{i}$ have a complicated dependence structure that cannot be reduced to independent random permutations.
@@
The aim of the remainder of this section is to sketch how the optimal spectral gap problem for the Laplacian $\Delta_{X^{N}}$ of a random cover is encoded as a strong convergence problem. This reduction proceeds in an analogous manner for the three models described above. We therefore fix in the following a base surface $\𝑋\ΓX=\Gamma\backslash\mathbb{H}$ and a sequence of random homomorphisms $\pi_{N}\in\mathrm{Hom}(\Gamma,\mathbf{S}_{N})$ as in any of the above models. The key assumption that will be needed, which holds in all three models, is that the random matrices $(U_{1}^{N},\ldots,U_{s}^{N})|_{1^{\perp}}$ defined by
@@
Beside models of random covers of hyperbolic surfaces, another important model of random surfaces is obtained by sampling from the Weil–Petersson measure on the moduli space of hyperbolic surfaces of genus $g$; this may be viewed as the natural notion of a typical surface of genus $g$. In a recent tour-de-force, Anantharaman and Monk [4, 5] proved that the Weil–Petersson model also exhibits an optimal spectral gap phenomenon by using methods inspired by Friedman’s original proof of Theorem 1.3. In contrast to random cover models, it does not appear that this problem can be reduced to a strong convergence problem. However, it is an interesting question whether a form of the polynomial method, which plays a key role in [92], could be used to obtain a new proof of this result.
@@ 5.2.2. Exploiting strong convergence
Recall that $e^{-\Delta_{\mathbb{H}}}$ is an integral operator on $L^{2}(\mathbb{H})$ with a smooth kernel $p_{\mathbb{H}}(x,y)$. The Laplacian $\Delta_{X^{N}}$ on $X^{N}=\Gamma\backslash(\mathbb{H}\times[N])$ is obtained by restricting the Laplacian on ×$\mathbb{H}\times[N]$ to functions that are invariant under $\Gamma$. In particular, this implies that $e^{-\Delta_{X^{N}}}$ may be viewed as an integral operator on $L^{2}(F\times[N])$ with kernel
[…]
by parameterizing $X^{N}$ as $F\times[N]$, where $F$ is the fundamental domain of the action of $\Gamma$ on $\mathbb{H}$. See, for example, [17, §3.7] or [70, §2].
@@ 2. $\bullet$
[…]
@@ 5.3. Random Schreier graphs
We may however ask what happens if we consider other actions of the symmetric group. Following [51, 30], denote by $[N]_{k}$ the set of all $k$-tuples of distinct elements of $[N]$. Then we obtain the natural action $\mathbf{S}_{N}\curvearrowright[N]_{k}$ by letting $\sigma$ act on each element of the tuple, that is, $\sigma(i_{1},\ldots,i_{k})=(\sigma(i_{1}),\ldots,\sigma(i_{k}))$. If we again choose $\sigma_{1},\ldots,\sigma_{r}$ to be i.i.d. uniform random elements of $\mathbf{S}_{N}$, then
@@ 2. $\bullet$
    The case $k=N$ corresponds to the Cayley graph of $\mathbf{S}_{N}$ with the random generators $\sigma_{1},\ldots,\sigma_{r}$, since $[N]_{N}\simeq\mathbf{S}_{N}$. Whether random Cayley graphs of $\mathbf{S}_{N}$ have an optimal spectral gap is a long-standing question (see section 6.2) that remains wide open: it has not even been shown that the maximum nontrivial eigenvalue is bounded away from the trivial eigenvalue in this setting.
@@
as $N\to\infty$ whenever $k_{N}\leq N^{\frac{1}{12}-\delta}$, for any $\delta>0$.
This yields a natural model of random $2r$-regular graphs with $|V|$ vertices that has an optimal spectral gap using only $\sim(\log|V|)^{12+\delta}$ bits of randomness, as compared to $\sim|V|\log|V|$ bits for ordinary random regular graphs.
@@
where $\pi_{N,k}:\mathbf{S}_{N}\to\mathrm{M}_{[N]_{k}}(\mathbb{C})$ maps $\sigma\in\mathbf{S}_{N}$ to the permutation matrix defined by its action on $[N]_{k}$. Then $\pi_{N,k}$ is clearly a group representation of $\mathbf{S}_{N}$, so it decomposes as a direct sum of irreducible representations $\pi_{N}^{\lambda}$. Theorem 5.7 now follows from the following result about strong convergence of irreducible representations of $\mathbf{S}_{N}$ that vastly generalizes Theorem 1.4 (which is the special case where $\pi_{N}^{\lambda}=\mathrm{std}_{N}$ is the standard representation, so that $\dim(\mathrm{std}_{N})=N-1$). For expository purposes, we state the result in a slightly more general form than is given in [30].
@@
$\lim_{N\to\infty}\max_{1<\dim(\pi_{N}^{\lambda})\leq\exp(N^{\frac{1}{12}-\delta})}\big{|}\|P(\pi_{N}^{\lambda}(\bm{\sigma}),\pi_{N}^{\lambda}(\bm{\sigma}^{*}))\|-\|P(\bm{u},\bm{u}^{*})\|\big{|}=0$
@@
The irreducible representations $\pi_{N}^{\lambda}$ are indexed by Young diagrams $\lambda\vdash N$. The argument in [47, §6] shows that for any $0<\delta^{\prime}<\delta$ and sufficiently large $N$, every irreducible representation with $\dim(\pi_{N}^{\lambda})\leq\exp(N^{\frac{1}{12}-\delta})$ has the property that the first row of either $\lambda$ or of the conjugate diagram $\lambda^{\prime}$ has length at least $N-N^{\frac{1}{12}-\delta^{\prime}}$. In the first case, the conclusion
$\lim_{N\to\infty}\max_{\lambda_{1}\geq N-N^{\frac{1}{12}-\delta^{\prime}}}\big{|}\|P(\pi_{N}^{\lambda}(\bm{\sigma}),\pi_{N}^{\lambda}(\bm{\sigma}^{*}))\|-\|P(\bm{u},\bm{u}^{*})\|\big{|}=0$ (5.1)
@@
On the other hand, as $\pi_{N}^{\lambda^{\prime}}(\sigma)=\mathrm{sgn}(\sigma)\pi_{N}^{\lambda}(\sigma)$ [78, Theorem 6.7], we obtain
$\lim_{N\to\infty}\max_{\lambda_{1}\geq N-N^{\frac{1}{12}-\delta^{\prime}}}\big{|}\|P(\pi_{N}^{\lambda^{\prime}}(\bm{\sigma}),\pi_{N}^{\lambda^{\prime}}(\bm{\sigma}^{*}))\|-\|P(\mathrm{sgn}(\bm{\sigma})\bm{u},\mathrm{sgn}(\bm{\sigma})\bm{u}^{*})\|\big{|}=0$
using that (5.1) holds uniformly over any finite set of polynomials $P$, and thus in particular over the polynomials $(\bm{u},\bm{u}^{*})\mapsto P(\bm{\varepsilon u},\bm{\varepsilon u}^{*})$ for all choices of signs $\bm{\varepsilon}\in\{-1,1\}^{r}$. It remains to note that $\|P(\mathrm{sgn}(\bm{\sigma})\bm{u},\mathrm{sgn}(\bm{\sigma})\bm{u}^{*})\|=\|P(\bm{u},\bm{u}^{*})\|$ by the Fell absorption principle [115, Proposition 8.1]. ∎
@@
Fix $\lambda\vdash k$, and consider the sequence of Young diagrams $\lambda(N)\vdash N$ (for $N\geq 2k$) so that removing the first row of $\lambda(N)$ yields $\lambda$; in particular, the first row has length $N-k$. Then the sequence of representations $\pi_{N}^{\lambda(N)}$ is called stable [48]. As was the case in section 3, any stable representation has the property that
@@
Taken more broadly, high dimensional representations of finite and matrix groups form a natural setting for the study of strong convergence and give rise to many interesting questions. For the unitary group $U(N)$, strong convergence was established earlier by Bordenave and Collins [21] for representations of polynomial dimension $N^{k}$, and by Magee and de la Salle [90] for representations of quasi-exponential dimension $\exp(N^{\alpha})$ (further improved in [33] using complementary ideas). On the other hand it is a folklore conjecture (see, e.g., [119, Conjecture 1.6]) that any sequence of representations of $\mathbf{S}_{N}$ of diverging dimension should give rise to optimal spectral gaps; Theorem 5.8 is at present the best known result in this direction. Analogous questions for finite simple groups of Lie type remain entirely open.

### v2

@@ 4.3. Discussion: on the role of interpolation
In the special case that $X$ is a GUE matrix, we obtained a much better result than Lemma 4.7: the crossing identity (4.1) captures the effect of a crossing exactly. This identity can be iterated in the presence of multiple crossings, which results in the genus expansion for GUE matrices (see, e.g., [101, §1.7]). This is a rather special feature of classical random matrix models, however, and we do not know of any method that can meaningfully capture the effect of multiple crossings in the setting of arbitrarily structured random matrices.
@@ 5.1. Random lifts of graphs
We begin by recalling some basic notions that can be found, for example, in [72, §6].
@@ 5.1.1. Random lifts
From the perspective of the lower bound, there is nothing particularly special about $d$-regular graphs beside that they all have the same universal cover. Indeed, for any sequence of graphs with diverging number of vertices that have the same universal cover, the maximum nontrivial eigenvalue is asymptotically lower bounded by the spectral radius of the universal cover. This follows by a straightforward adaptation of Lemma 1.2, cf. [72, Theorem 6.6].
@@ 5.1.2. Old and new eigenvalues
Even though Corollary 5.2 is only concerned with the new eigenvalues of $G^{N}$, it implies the classical spectral gap property $\|A^{N}|_{1^{\perp}}\|\to\varrho$ whenever the base graph satisfies $\|A|_{1^{\perp}}\|\leq\varrho$. Another simple consequence is that whenever the base graph satisfies $\|A\|>\varrho$, the random lift $G^{N}$ is connected with probability 1o(1)1-o(1); this holds if and only if $G$ has at least two cycles [74, Theorem 2].
@@ 5.2. Buser’s conjecture
[…]
@@ 5.2.1. Random covers
[…]
$F\scriptstyle\gamma_{1}F$$\scriptstyle\gamma_{2}^{\text{-}1}F$$\scriptstyle\gamma_{1}^{\text{-}1}F$$\scriptstyle\gamma_{2}F$$\scriptstyle\gamma_{3}F$$\scriptstyle\gamma_{4}^{\text{-}1}F$$\scriptstyle\gamma_{3}^{\text{-}1}F$$\scriptstyle\gamma_{4}F$$\Gamma\backslash\mathbb{H}$
Figure 5.2. Illustration of a tiling of $\mathbb{H}$ (in the Poincaré disc model) by hyperbolic octagons; gluing the sides of the fundamental domain $F$ yields a genus 22 surface.
@@
[…]
@@
where we let $\gamma\in\Gamma$ act on $\mathbb{H}\times[N]$ as $\gamma(z,i)=(\gamma z,\pi_{N}(\gamma)i)$. Then $X^{N}$ is an $N$-fold cover of $X$, and every $N$-fold cover of $X$ arises in this manner for some choice of $\pi_{N}$; cf. [64, pp. 68–70] or [53, §14a and §16d].
@@
where $g,h]=ghg^{-1}h^{-1}$. In this case, the random permutations $\sigma_{i}$ must be chosen to satisfy $\sigma_{1},\sigma_{2}]\cdots[\sigma_{2g-1},\sigma_{2g}]=1$, which precludes them from being independent. The reason this issue does not arise for graphs is that the fundamental group of every graph is free, and thus there are no relations to be satisfied.
@@
    Finally, the most natural model of random covers of closed surfaces is to choose $\pi_{N}\in\mathrm{Hom}(\Gamma,\mathbf{S}_{N})$ uniformly at random, that is, choose $\sigma_{i}=\pi_{N}(\gamma_{i})$ uniformly at random among the set of tuples $\sigma_{1},\ldots,\sigma_{s}\in\mathbf{S}_{N}$ that satisfy the relation of $\Gamma$. This corresponds to choosing an $N$-fold cover of $X$ uniformly at random [92, 93]. The challenge in analyzing this model is that $\sigma_{i}$ have a complicated dependence structure that cannot be reduced to independent random permutations.
@@
[…]
@@
Beside models of random covers of hyperbolic surfaces, another important model of random surfaces is obtained by sampling from the Weil–Petersson measure on the moduli space of hyperbolic surfaces of genus $g$; this may be viewed as the natural notion of a typical surface of genus $g$. In a tour-de-force, Anantharaman and Monk [4, 5] proved that the Weil–Petersson model also exhibits an optimal spectral gap phenomenon by using methods inspired by Friedman’s original proof of Theorem 1.3. In contrast to random cover models, it does not appear that this result can be reduced to a strong convergence problem. Nonetheless, it was recently shown by Hide–Macera–Thomas [69] that the the polynomial method, which plays a key role in [94], can be applied directly to achieve a new proof of this result.
@@ 5.2.2. Exploiting strong convergence
Recall that $e^{-\Delta_{\mathbb{H}}}$ is an integral operator on $L^{2}(\mathbb{H})$ with a smooth kernel $p_{\mathbb{H}}(x,y)$. The Laplacian $\Delta_{X^{N}}$ on $X^{N}=\Gamma\backslash(\mathbb{H}\times[N])$ is obtained by restricting the Laplacian on $\mathbb{H}\times[N]$ to functions that are invariant under $\Gamma$. In particular, this implies that $e^{-\Delta_{X^{N}}}$ may be viewed as an integral operator on $L^{2}(F\times[N])$ with kernel
$p_{X^{N}}((x,i),(y,j))=\sum_{\gamma\in\Gamma}p_{\mathbb{H}}(x,\gamma y)\,1_{i=\pi_{N}(\gamma)j}$
by parameterizing $X^{N}$ as $F\times[N]$, where $F$ is the fundamental domain of the action of $\Gamma$ on $\mathbb{H}$. See, for example, [17, §3.7] or [71, §2].
@@ 2. $\bullet$
$b\tilde{f}(x,g)=\sum_{\gamma\in\Gamma}\int_{F}p_{\mathbb{H}}(x,\gamma y)\,f(g^{-1}\gamma y)\,dy=\int_{\mathbb{H}}p_{\mathbb{H}}(g^{-1}x,y)\,f(y)\,dy=\widetilde{e^{-\Delta_{\mathbb{H}}}f}(x,g),$
@@ 5.3. Random Schreier graphs
We may however ask what happens if we consider other actions of the symmetric group. Following [51, 30], denote by $N]_{k}$ the set of all $k$-tuples of distinct elements of $[N]$. Then we obtain the natural action $\mathbf{S}_{N}\curvearrowright[N]_{k}$ by letting $\sigma$ act on each element of the tuple, that is, $\sigma(i_{1},\ldots,i_{k})=(\sigma(i_{1}),\ldots,\sigma(i_{k}))$. If we again choose $\sigma_{1},\ldots,\sigma_{r}$ to be i.i.d. uniform random elements of $\mathbf{S}_{N}$, then
@@ 2. $\bullet$
    The case $k=N$ corresponds to the Cayley graph of $\mathbf{S}_{N}$ with the random generators $\sigma_{1},\ldots,\sigma_{r}$, since $N]_{N}\simeq\mathbf{S}_{N}$. Whether random Cayley graphs of $\mathbf{S}_{N}$ have an optimal spectral gap is a long-standing question (see section 6.2) that remains wide open: it has not even been shown that the maximum nontrivial eigenvalue is bounded away from the trivial eigenvalue in this setting.
@@
as $N\to\infty$ whenever $k_{N}\leq N^{\frac{1}{20}-\delta}$, for any $\delta>0$.
This yields a natural model of random $2r$-regular graphs with $|V|$ vertices that has an optimal spectral gap using only $\sim(\log|V|)^{20+\delta}$ bits of randomness, as compared to $\sim|V|\log|V|$ bits for ordinary random regular graphs.
@@
where $\pi_{N,k}:\mathbf{S}_{N}\to\mathrm{M}_{[N]_{k}}(\mathbb{C})$ maps $\sigma\in\mathbf{S}_{N}$ to the permutation matrix defined by its action on $N]_{k}$. Then $\pi_{N,k}$ is clearly a group representation of $\mathbf{S}_{N}$, so it decomposes as a direct sum of irreducible representations $\pi_{N}^{\lambda}$. Theorem 5.7 now follows from the following result about strong convergence of irreducible representations of $\mathbf{S}_{N}$ that vastly generalizes Theorem 1.4 (which is the special case where $\pi_{N}^{\lambda}=\mathrm{std}_{N}$ is the standard representation, so that $\dim(\mathrm{std}_{N})=N-1$). For expository purposes, we state the result in a slightly more general form than is given in [30].
@@
$\lim_{N\to\infty}\max_{1<\dim(\pi_{N}^{\lambda})\leq\exp(N^{\frac{1}{20}-\delta})}\big|\|P(\pi_{N}^{\lambda}(\bm{\sigma}),\pi_{N}^{\lambda}(\bm{\sigma}^{*}))\|-\|P(\bm{u},\bm{u}^{*})\|\big|=0$
@@
The irreducible representations $\pi_{N}^{\lambda}$ are indexed by Young diagrams $\lambda\vdash N$. The argument in [47, §6] shows that for any $0<\delta^{\prime}<\delta$ and sufficiently large $N$, every irreducible representation with $\dim(\pi_{N}^{\lambda})\leq\exp(N^{\frac{1}{20}-\delta})$ has the property that the first row of either $\lambda$ or of the conjugate diagram $\lambda^{\prime}$ has length at least $N-N^{\frac{1}{20}-\delta^{\prime}}$. In the first case, the conclusion
$\lim_{N\to\infty}\max_{\lambda_{1}\geq N-N^{\frac{1}{20}-\delta^{\prime}}}\big|\|P(\pi_{N}^{\lambda}(\bm{\sigma}),\pi_{N}^{\lambda}(\bm{\sigma}^{*}))\|-\|P(\bm{u},\bm{u}^{*})\|\big|=0$ (5.1)
@@
On the other hand, as $\pi_{N}^{\lambda^{\prime}}(\sigma)=\mathrm{sgn}(\sigma)\pi_{N}^{\lambda}(\sigma)$ [79, Theorem 6.7], we obtain
$\lim_{N\to\infty}\max_{\lambda_{1}\geq N-N^{\frac{1}{20}-\delta^{\prime}}}\big|\|P(\pi_{N}^{\lambda^{\prime}}(\bm{\sigma}),\pi_{N}^{\lambda^{\prime}}(\bm{\sigma}^{*}))\|-\|P(\mathrm{sgn}(\bm{\sigma})\bm{u},\mathrm{sgn}(\bm{\sigma})\bm{u}^{*})\|\big|=0$
using that (5.1) holds uniformly over any finite set of polynomials $P$, and thus in particular over the polynomials $(\bm{u},\bm{u}^{*})\mapsto P(\bm{\varepsilon u},\bm{\varepsilon u}^{*})$ for all choices of signs $\bm{\varepsilon}\in\{-1,1\}^{r}$. It remains to note that $\|P(\mathrm{sgn}(\bm{\sigma})\bm{u},\mathrm{sgn}(\bm{\sigma})\bm{u}^{*})\|=\|P(\bm{u},\bm{u}^{*})\|$ by the Fell absorption principle [117, Proposition 8.1]. ∎
@@
Fix $\lambda\vdash k$, and consider the sequence of Young diagrams $\lambda(N)\vdash N$ (for $N\geq 2k$) so that removing the first row of $\lambda(N)$ yields $\lambda$; in particular, the first row has length NkN-k. Then the sequence of representations $\pi_{N}^{\lambda(N)}$ is called stable [48]. As was the case in section 3, any stable representation has the property that
@@
Taken more broadly, high dimensional representations of finite and matrix groups form a natural setting for the study of strong convergence and give rise to many interesting questions. For the unitary group $U(N)$, strong convergence was established earlier by Bordenave and Collins [21] for representations of polynomial dimension $N^{k}$, and by Magee and de la Salle [91] for representations of quasi-exponential dimension $\exp(N^{\alpha})$ (further improved in [33] using complementary ideas). On the other hand it is a folklore conjecture (see, e.g., [121, Conjecture 1.6]) that any sequence of representations of $\mathbf{S}_{N}$ of diverging dimension should give rise to optimal spectral gaps; Theorem 5.8 is at present the best known result in this direction. Analogous questions for finite simple groups of Lie type remain entirely open.